OpenStudy (anonymous):

a population grows exponentially and doubles in size every 10 years. using P=P0e^kt, where Po represents the population at t=0, find the value of k

OpenStudy (lgbasallote):

here's a shortcut you can learn \[k = \frac{\ln 2}{t}\] where t is the time where the population doubles in size so \[k = \frac{\ln 2}{10}\] does that help?

OpenStudy (anonymous):

doubles in ten years so you know \[2=e^{k\times 10}\] solve for \(k\) yo get \[\ln(2)=10k\] and so \[k=\frac{\ln(2)}{10}\] just as written above